how many pounds of an alloy of 20% silver must be melted with 50lbs of an alloy with 30% silver in order to get an alloy of 27% silver
Answer by checkley77(12844)

.20x+.30*50=.27(50+x)
.20x+15=13.5+.27x
.20x-.27x=13.5-15
-.07x=-1.5
x=-1.5/-.07
x=21.43 pounds of 20% silver is used.
Proof:
.20*21.43+.30*50=.27(50+21.43)
4.286+15=.27*71.43
19.286=19.286